# X-Troll: eXplainable Detection of State-Sponsored Information Operations Agents

### v1

@@ Abstract.
State-sponsored trolls, malicious actors who deploy sophisticated linguistic manipulation in coordinated information campaigns, posing threats to online discourse integrity. While Large Language Models (LLMs) achieve strong performance on general natural language processing (NLP) tasks, they struggle with subtle propaganda detection and operate as “black boxes”, providing no interpretable insights into manipulation strategies. This paper introduces X-Troll, a novel framework that bridges this gap by integrating explainable adapter-based LLMs with expert-derived linguistic knowledge to detect state-sponsored trolls and provide human-readable explanations for its decisions. X-Troll incorporates appraisal theory and propaganda analysis through specialized LoRA adapters, using dynamic gating to capture campaign-specific discourse patterns in coordinated information operations. Experiments on real-world data demonstrate that our linguistically-informed approach shows strong performance compared with both general LLM baselines and existing troll detection models in accuracy while providing enhanced transparency through expert-grounded explanations that reveal the specific linguistic strategies used by state-sponsored actors. X-Troll source code is available at: https://anonymous.4open.science/r/xtroll_source-6656/.
@@ 1. Introduction
In this paper, we study the automated detection of state-sponsored troll users based on their social media post timelines, formulating the problem as a linguistically-informed text classification task that goes beyond general LLM capabilities. To address the limitations revealed by recent LLM evaluations and leverage the insights from linguistic analysis, we pose two key research questions:
@@ 4.1. Problem Statement
Rationale Selection and Explanation Generation. We address explainability through a two-stage interpretability mechanism. First, a rationale selector $f_{\gamma}$ identifies a sparse subset of $k$ salient tokens across the user’s timeline. These tokens most strongly influence the classification decision: $\mathcal{R}_{u}=f_{\gamma}(T_{u},y_{u})$. Second, an explanation generator $s_{\omega}$ synthesizes a comprehensive, human-readable explanation from the extracted rationales: $S_{u}=s_{\omega}(\mathcal{R}_{u})$. This explanation articulates the specific behavioral and linguistic patterns that characterize the troll activity.
@@ 5.1. Datasets
For troll detection and information campaign classification tasks, we used datasets from three specific state-sponsored information campaigns: Russia-Anti-NATO Troll, Russia-IRA Troll, and PRC-Xinjiang Troll [^3] https://blog.x.com/en_us/topics/company/2021/disclosing-state-linked-information-operations-we-ve-removed. These datasets were released by Twitter and contain troll accounts and their posts. Twitter banned these accounts in October 2018 and anonymized them to ensure they could not be associated with individual users. For example, the ‘Russia-Anti-NATO Troll’ dataset contains users banned for amplifying narratives that sought to undermine faith in the NATO alliance and its stability [^4] https://blog.x.com/en_us/topics/company/2021/disclosing-networks-of-state-linked-information-operations-. We also collected posts from non-troll users via the Twitter data API to establish a baseline dataset for evaluating model performance on non-troll data. In our non-troll data collection, we excluded all profile information and retrieved only necessary data. We focused on posts related to specific topics to protect user privacy. Table 1 presents the summary of the data statistics including the number of users and posts. To enhance analysis quality, domain experts annotated a subset of posts with appraisal labels for each information campaign, as shown in Table 1. The propaganda strategy adapter was fine-tuned on datasets provided by Moral et al. (2023) as part of the DIPROMATS 2023 shared task [^5] https://sites.google.com/view/dipromats2023.
@@ 5.4. Campaign-Specific Patterns
Figure 3. Rationale selector in X-Troll applied to the Russia-IRA dataset. Rationale words are highlighted. Top: a correctly identified troll user. Bottom: a misclassified non-troll user.
@@ 5.5. Explainability Analysis
Summary Generation Evaluation. Table 3 reveals certain patterns in explanation generation performance. The rationale selector is able to improve explanation quality across most model-dataset combinations, with strong improvements for FLAN-T5 on Russia-IRA data (7.17.1% improvement) and Falcon-7B on PRC-Xinjiang data (6.26.2% improvement). However, some configurations show slight degradation (FLAN-T5 on Russia-Anti-NATO: 4.5-4.5%), suggesting that rationale effectiveness varies with campaign characteristics and base model capabilities. We provide detailed evaluation prompts and templates used for G-Eval scoring (Liu et al., 2023b) in Appendix B.
@@ Acknowledgements
Appendix A Summary Evaluation
@@
Appendix B Summary Evaluation Prompt Template

### v2

@@ Abstract.
State-sponsored trolls, malicious actors who deploy sophisticated linguistic manipulation in coordinated information campaigns, posing threats to online discourse integrity. While Large Language Models (LLMs) achieve strong performance on general natural language processing (NLP) tasks, they struggle with subtle propaganda detection and operate as “black boxes”, providing no interpretable insights into manipulation strategies. This paper introduces X-Troll, a novel framework that bridges this gap by integrating explainable adapter-based LLMs with expert-derived linguistic knowledge to detect state-sponsored trolls and provide human-readable explanations for its decisions. X-Troll incorporates appraisal theory and propaganda analysis through specialized LoRA adapters, using dynamic gating to capture campaign-specific discourse patterns in coordinated information operations. Experiments on real-world data demonstrate that our linguistically-informed approach shows strong performance compared with both general LLM baselines and existing troll detection models in accuracy while providing enhanced transparency through expert-grounded explanations that reveal the specific linguistic strategies used by state-sponsored actors. X-Troll source code is available at: https://github.com/ltian678/xtroll_source/.
@@ 1. Introduction
To address the limitations revealed by recent LLM evaluations and leverage the insights from linguistic analysis, we pose two key research questions:
@@ 4.1. Problem Statement
Rationale Selection and Explanation Generation. We address explainability through a two-stage interpretability mechanism. First, a rationale selector $f_{\gamma}$ identifies a sparse subset of $k$ salient tokens across the user’s timeline. These tokens most strongly influence the classification decision: $\mathcal{R}_{u}=f_{\gamma}(T_{u},y_{u})$. Second, an explanation generator $s_{\omega}$ synthesizes a comprehensive, human-readable explanation from the extracted rationales: $S_{u}=s_{\omega}(\mathcal{R}_{u})$.
@@ 5.1. Datasets
For troll detection and information campaign classification tasks, we used datasets from three specific state-sponsored information campaigns: Russia-Anti-NATO Troll, Russia-IRA Troll, and PRC-Xinjiang Troll [^3] https://blog.x.com/en_us/topics/company/2021/disclosing-state-linked-information-operations-we-ve-removed. These datasets were released by Twitter and contain troll accounts and their posts. Twitter banned these accounts in October 2018 and anonymized them to ensure they could not be associated with individual users. For example, the ‘Russia-Anti-NATO Troll’ dataset contains users banned for amplifying narratives that sought to undermine faith in the NATO alliance and its stability [^4] https://blog.x.com/en_us/topics/company/2021/disclosing-networks-of-state-linked-information-operations-. We also collected posts from non-troll users via the Twitter data API to establish a baseline dataset for evaluating model performance on non-troll data. In our non-troll data collection, we excluded all profile information and retrieved only necessary data. We focused on posts related to specific topics to protect user privacy. Table 1 presents the summary of the data statistics including the number of users and posts. To enhance analysis quality, domain experts annotated a subset of posts with appraisal labels for each information campaign, as shown in Table 1. The propaganda strategy adapter was fine-tuned on datasets provided by Moral et al. (2023) as part of the DIPROMATS 2023 shared task [^5] https://sites.google.com/view/dipromats2023. We also provide a sample of annotated posts with detailed labels (including appraisal and propaganda tags), which is included in Appendix (app, ndix).
@@ 5.4. Campaign-Specific Patterns
Figure 3. X-Troll’s rationale selection on Russia-IRA examples, with diagnostic tokens highlighted. The correctly classified troll post (top) shows characteristic geopolitical framing and conflict narratives, while the false positive (bottom) reveals how political topic overlap without coordinated rhetorical patterns can mislead classification.
@@ 5.5. Explainability Analysis
Summary Generation Evaluation. Table 3 reveals certain patterns in explanation generation performance. The rationale selector is able to improve explanation quality across most model-dataset combinations, with strong improvements for FLAN-T5 on Russia-IRA data (7.17.1% improvement) and Falcon-7B on PRC-Xinjiang data (6.26.2% improvement). However, some configurations show slight degradation (FLAN-T5 on Russia-Anti-NATO: 4.5-4.5%), suggesting that rationale effectiveness varies with campaign characteristics and base model capabilities. In Appendix (app, ndix), we include the detailed G-Eval prompts and templates used for scoring (Liu et al., 2023b), as well as the complete evaluation results, reporting the scores for each evaluation criterion.
@@ Acknowledgements
Appendix A Sample Annotated Data
We provide a sample of our annotated data with detailed labels (see Fig. 5). While the main experiments rely on large-scale datasets described in Section 5.1, this illustrative sample shows our annotation schema and the kinds of linguistic patterns X-Troll is designed to capture.
State-sponsored troll detection requires moving beyond post-level classification to understand how manipulation strategies unfold linguistically across different narrative contexts. Our annotation framework captures the complex interplay between ideational targeting, evaluative stance, and persona construction that characterizes coordinated inauthentic behavior.
Our annotation framework captures three interconnected dimensions of state-sponsored discourse. The ideational target dimension identifies which entities—from specific groups—are being framed in each post, revealing strategic attention direction. The appraisal dimension encodes evaluative stance toward these targets, capturing subtle positioning shifts that traditional sentiment analysis might miss. The persona dimension characterizes the discursive role adopted by accounts, showing how agents modulate voice for credibility and detection evasion.
The sample shows 2020 Syrian conflict discourse, where these dimensions interact to produce sophisticated manipulation. Posts combining “reportage” personas with selective military targeting exemplify how neutral news-sharing masks geopolitical agenda-setting.
Figure 5. Sample of annotated posts from the Russia–Syria dataset, showing text excerpts with corresponding expert labels (ideational target, persona, appraisal, and propaganda technique). Full datasets are described in Section 5.1.
Appendix B Summary Evaluation
@@
Appendix C Summary Evaluation Prompt Template